### v1

# Model-free source seeking of exponentially convergent unicycle: theoretical and robotic experimental results

Abstract
This paper introduces a novel model-free, real-time unicycle-based source seeking design. This design steers autonomously the unicycle dynamic system towards the extremum point of an objective function or physical/scaler signal that is unknown expression-wise, but accessible via measurements. A key contribution of this paper is that the introduced design converges exponentially to the extremum point of objective functions (or scaler signals) that behave locally like a higher-degree power functions (e.g., fourth degree polynomial function) as opposed to locally quadratic objective functions, the usual case in literature. We provide theoretical and simulation results to support out theoretical results. Also, for the first time in the literature, we provide experimental robotic results that demonstrate the effectiveness of the proposed design and its exponential convergence ability.
keywords:
Source Seeking, Extremum Seeking, Unicycle, Exponential Convergence, Robotic Experiment, Model-free Optimization, Light source seeking.
1 Introduction
Extremum seeking control (ESC) techniques are model-free, real-time control methods that drive a system to the optimum of a given objective function (Ariyur and Krstic (2003); Tan et al. (2010); Guay and Dochain (2015); Scheinker (2024)), which may be unknown, provided it can be measured. ESC schemes are particularly attractive across various fields due to their minimal information requirements: they rely solely on applying perturbations to system parameters or inputs, and measuring the corresponding outputs of the objective function. Using these measurements in a feedback loop, ESC algorithms iteratively adjust the system input or parameter values to drive it towards the extremum, as can be seen in many applications, e.g., Krstic and Cochran (2008); Dochain et al. (2011); Calli et al. (2012); Bajpai et al. (2024); Elgohary et al. (2025); Elgohary and Eisa (2025a, b); Pokhrel and Eisa (2022); Eisa and Pokhrel (2023); Grushkovskaya et al. (2018a); Grushkovskaya and Zuyev (2024).
One conventional problem that has increasingly involved the use of ESC methods is source-seeking (Bajpai (2024); Ghods (2011); Zhu et al. (2013); Li et al. (2014); Bulgur et al. (2018)). Source-seeking is the problem of steering a system (e.g., a mobile robot) autonomously to the maximum or minimum intensity of a physical/scaler signal present in a given domain using only on-board sensor measurements (e.g., heat, light, chemical concentration, among others). The advantage of ESC lies in the fact that it does not require knowledge of the signal’s distribution or explicit modeling of the system to drive the system, autonomously, to the extremum point. Due to the model-free nature of ESC methods, source-seeking techniques based on ESC utilized a simple, basic unicycle model to represent the system within the control design (Pokhrel et al. (2024); Matveev et al. (2011); Khong et al. (2014); Cochran et al. (2009); Grushkovskaya et al. (2018a); Suttner and Krstić (2022); Yilmaz et al. (2025); Todorovski and Krstić (2024); Elgohary et al. (2025)). As a result, unicycle-based methods for source-seeking often resort to Lie bracket averaging (Dürr et al. (2013); Scheinker (2017); Grushkovskaya et al. (2018b); Ghadiri-Modarres and Mojiri (2020); Pokhrel and Eisa (2023)) in theoretical analysis and design.
Motivation and Contribution. In our recent work Grushkovskaya and Eisa (2025), new ESC laws were derived based on higher-order Lie bracket averaging. This paved the way for the prospect of having exponentially convergent ESC laws for objective functions that behave locally not as quadratic functions, but as power functions with higher-order degrees. Inspired by the prospect of realizing exponential convergence for high-order objective functions, we seek to utilize the ESC method in Grushkovskaya and Eisa (2025) and propose a novel, first-of-its-kind, exponentially convergent unicycle-based design for source-seeking. In this preliminary study, we only focus on source-seeking of objective functions that behave locally as a fourth-degree power function. We prove stability of the proposed design and provide both simulation and experimental results. Our experimental demonstration uses TurtleBot3 robots (similar to Bajpai et al. (2024); Elgohary et al. (2025)) and we compare our design (third-order Lie bracket-based) with traditional results using first-order Lie bracket-based design. We also perform a light source-seeking experiment in a complete model-free fashion. Our results illustrate the potential and effectiveness of the proposed exponentially convergent unicycle design for source-seeking.
2 Brief Background
In this section, we provide brief context regarding the exponentially convergent ESC law that we use for high-order objective functions. We consider the two-input control-affine ESC of the following form as in Grushkovskaya and Eisa (2025):
$\dot{x}=g_{1}(J(x))u_{1}^{\varepsilon}(t)+g_{2}(J(x))u_{2}^{\varepsilon}(t),$ (1)
where $u_{1}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{1}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{2}(t/\varepsilon)$, $\varepsilon>0$, such that the dithers, $v_{1}(t/\varepsilon)$ and $v_{2}(t/\varepsilon)$ excite the Lie bracket $g_{I_{N}}(z)=\big[\big[\dots[g_{1},\underbrace{g_{2}],g_{2}\big],\dots,g_{2}\big]}_{N-1\ \text{times}}\dots]\big](z)$ at time $t=\varepsilon$ with $I_{N}=(1,\underbrace{2,\dots,2}_{N-1\ \text{times}})$, where
$g_{i},g_{j}]:=\frac{\partial g_{j}}{\partial x}g_{i}-\frac{\partial g_{i}}{\partial x}g_{j}.$
Moreover, we adopt all mathematical assumptions on the vector fields $g_{1}$, $g_{2}$ provided in (Grushkovskaya and Eisa, 2025, Section II). The selection of the dither signals is done as in Grushkovskaya and Zuyev (2024); Grushkovskaya and Eisa (2025), conditions C1-C3:
  • C1. With, $v_{1}(t/\varepsilon)=2\sqrt{\kappa_{12}\pi}\cos{(2\kappa_{12}\pi t/\varepsilon)}$ and $v_{2}(t/\varepsilon)=2\sqrt{\kappa_{12}\pi}\sin{(2\kappa_{12}\pi t/\varepsilon)}$, $\kappa_{12}\in\mathbb{Z}$, the first-order Lie bracket is excited.
  • C2. With, $v_{1}(t/\varepsilon)=-2(4\kappa_{122}\pi)^{2/3}\cos{(4\kappa_{122}\pi t/\varepsilon)}$ and $v_{2}(t/\varepsilon)=(4\kappa_{122}\pi)^{2/3}\cos{(2\kappa_{122}\pi t/\varepsilon)}$, $\kappa_{122}\in\mathbb{Z}$, the second-order Lie bracket is excited.
  • C3. With, $v_{1}(t/\varepsilon)=6(2\kappa_{1222}\pi)^{3/4}\sin{(6\kappa_{1222}\pi t/\varepsilon)}$ and $v_{2}(t/\varepsilon)=2(2\kappa_{1222}\pi)^{3/4}\cos{(2\kappa_{1222}\pi t/\varepsilon)}$, $\kappa_{1222}\in\mathbb{Z}$, the third-order Lie bracket is excited.
From Grushkovskaya and Eisa (2025), $g_{1},\ g_{2}$ are chosen according to:
$g_{I_{N}}(J(x))=-c_{N}J^{(N-1)}(x),$ (2)
with the simple selection to satisfy (2) being $g_{1}(z)=-1^{(N+1)}z$ and $g_{2}(z)=1$. It is also to be noted that the conditions C1-C3 above guarantee that only the desired Lie bracket is excited while all others vanish per Gauthier and Kawski (2014); Pokhrel and Eisa (2023); Grushkovskaya and Zuyev (2024). In this paper, we will be interested in third-order Lie bracket excitation (i.e., condition C3).
3 Main Results: Proposed Design and Stability Analysis
We now propose our exponentially convergent unicycle design and prove its stability.
3.1 Proposed Design
We aim at using an ESC law based on the conditions C1-C3 provided in the previous section to achieve an exponentially convergent unicycle design. Let us consider the kinematic differential equations for unicycle dynamics with constant angular velocity as follows (Elgohary et al. (2025); Grushkovskaya et al. (2018a)):
$\dot{x}$ $=\mathrm{v}\cos(\Omega t)$ (3)
$\dot{y}$ $=\mathrm{v}\sin(\Omega t),$
where $x$ and $y$ describe the current position/coordinates, $\mathrm{v}$ is the linear/transnational velocity, and $\Omega$ is the angular velocity. While the results can be extended to any order, in this study, we only focus the unicycle design to address objective functions that behave locally as fourth-order degree polynomial in consistency with (Grushkovskaya and Eisa, 2025, Assumption 1). Hence, we suppose:
$J(x,y)=C_{1}(x-x_{d})^{4}+C_{2}(y-y_{d})^{4},$ (4)
where $C_{1},C_{2}>0$. As analyzed, shown and simulated in Grushkovskaya and Eisa (2025), we need to excite third-order Lie bracket for exponential convergence since the objective function order is a fourth-degree. Hence, we choose our control law for $\mathrm{v}$ of the unicycle based on condition C3. That is,
$\mathrm{v}=2(2\pi/\varepsilon)^{(3/4)}(3cJ(x,y)\sin{(6\pi t/\varepsilon)}+a\cos{(2\pi t/\varepsilon)}),$ (5)
with $c,a>0$. Now we are in a position to put all elements of the proposed design together, which is depicted in Figure 1. In state space representation, including an optional high-pass filter (HPF) into the system (Bajpai et al. (2024); Elgohary et al. (2025)), the proposed design becomes:
$\dot{x}$ $=\big(2(2\pi/\varepsilon)^{(3/4)}(3c(J-eh)\sin{(6\pi t/\varepsilon)}+a\cos{(2\pi t/\varepsilon)})\big)\cos{(\Omega t)},$ (6)
$\dot{y}$ $=\big(2(2\pi/\varepsilon)^{(3/4)}(3c(J-eh)\sin{(6\pi t/\varepsilon)}+a\cos{(2\pi t/\varepsilon)})\big)\sin{(\Omega t)},$
$\dot{h}$ $=J(x,y)-eh,$
where $h$ is the optional filter state and $e$ is the filter constant. If the design is to be considered without HPF, we set $h=0$ and omit its dynamical equation.
Figure 1: The proposed ESC design for exponentially convergent unicycle. We use the proposed unicycle design for differential drive robotic experiments. In experiments and simulations we use mathematically known objective function and unknown light source.
3.2 Stability Analysis
In this subsection, we provide the stability analysis for the proposed unicycle design (6). We will follow the traditional methodology in Lie bracket-based ESC literature (e.g., Dürr et al. (2013); Grushkovskaya et al. (2018b); Pokhrel and Eisa (2023)) where the stability property of the ESC system is characterized by the corresponding Lie bracket system (LBS). The corresponding LBS to (6) without the optional HPF (i.e., $h=0$) is a third-order LBS based on condition C3 of the form (Grushkovskaya and Eisa (2025)):
$\dot{\bar{x}}=\big[[[g_{1},g_{2}],g_{2}],g_{2}\big]$
with
$g_{1}=\left(\begin{array}[]{c}cJ(\bar{x},\bar{y})\cos(\Omega t)\\ cJ(\bar{x},\bar{y})\sin(\Omega t)\end{array}\right),\,g_{2}=\left(\begin{array}[]{c}a\cos(\Omega t)\\ a\sin(\Omega t)\end{array}\right).$
Let $J_{\bar{x}\bar{x}\bar{x}}$ denotes the third order partial derivative of $J(\bar{x},\bar{y})$ with respect to $\bar{x}$ and $J_{\bar{y}\bar{y}\bar{y}}$ denotes the third order partial derivative of $J(\bar{x},\bar{y})$ with respect to $\bar{y}$. Then,
$\dot{\bar{x}}$ $=-ca^{3}\cos(\Omega t)(J_{\bar{x}\bar{x}\bar{x}}\cos^{3}(\Omega t)+J_{\bar{y}\bar{y}\bar{y}}\sin^{3}(\Omega t))$ (7)
$=-\cos(\Omega t)(c_{1}(\bar{x}-\bar{x}_{d})\cos^{3}(\Omega t)+c_{2}(\bar{y}-\bar{y}_{d})\sin^{3}(\Omega t)),$
$\dot{\bar{y}}$ $=-ca^{3}\sin(\Omega t)(J_{\bar{x}\bar{x}\bar{x}}\cos^{3}(\Omega t)+J_{\bar{y}\bar{y}\bar{y}}\sin^{3}(\Omega t))$
$=-\sin(\Omega t)(c_{1}(\bar{x}-\bar{x}_{d})\cos^{3}(\Omega t)+c_{2}(\bar{y}-\bar{y}_{d})\sin^{3}(\Omega t)),$
with $c_{1}=4!ca^{3}C_{1}$, $c_{2}=4!ca^{3}C_{2}$. The following result establishes the stability properties of system (7).
Theorem 1
Let one of the following conditions be satisfied:
  • i)
    $c_{1}\in(3c_{2}/5,c_{2}]$, $\Omega>\frac{2c_{1}(3c_{2}-c_{1})(2c_{2}-c_{1})}{16c_{1}^{2}-(3c_{2}-c_{1})^{2}}$;
  • ii)
    $c_{2}\in(3c_{1}/5,c_{1}]$, $\Omega>\frac{2c_{2}(3c_{1}-c_{2})(2c_{1}-c_{2})}{16c_{2}^{2}-(3c_{1}-c_{2})^{2}}$.
Then the equilibrium $x^{*}=x_{d},y^{*}=y_{d}$ of system (7) is exponentially stable.
Proof. We begin with the following change of variables:
$\xi=(\bar{x}-\bar{x}_{d})\cos(\Omega t)+(\bar{y}-\bar{y}_{d})\sin(\Omega t),$
$\eta=(\bar{x}-\bar{x}_{d})\sin(\Omega t)-(\bar{y}-\bar{y}_{d})\cos(\Omega t),$
with the inverse transformation
$\bar{x}-\bar{x}_{d}=\xi\cos(\Omega t)+\eta\sin(\Omega t),$
$\bar{y}-\bar{y}_{d}=\xi\sin(\Omega t)-\eta\cos(\Omega t).$
Observe that
$\dot{\xi}=$ $\dot{\bar{x}}\cos(\Omega t)+\dot{\bar{y}}\sin(\Omega t)$
$-\Omega((\bar{x}-\bar{x}_{d})\sin(\Omega t)-(\bar{y}-\bar{y}_{d})\cos(\Omega t))$
$=$ $-(c_{1}(x-x_{d})\cos^{3}(\Omega t)+c_{2}(y-y_{d})\sin^{3}(\Omega t))-\Omega\eta$
$=$ $-\xi(c_{1}\cos^{4}(\Omega t)+c_{2}\sin^{4}(\Omega t))$
$-\frac{1}{2}\eta\sin(2\Omega t)(c_{1}\cos^{2}(\Omega t)-c_{2}\sin^{2}(\Omega t)),$
and
$\dot{\eta}=$ $\dot{x}\sin(\Omega t)-\dot{y}\cos(\Omega t)$
$+\Omega((x-x_{d})\cos(\Omega t)+(y-y_{d})\sin(\Omega t))$
$=$ $\Omega\xi.$
Thus, in the new variables system (7) takes the form
$\dot{\xi}=-\kappa_{1}(t)\xi-\eta(\Omega+\kappa_{2}(t)),$ (8)
$\dot{\eta}=\Omega\xi.$
where
$\kappa_{1}(t)=c_{1}\cos^{4}(\Omega t)+c_{2}\sin^{4}(\Omega t),$
$\kappa_{2}(t)=\frac{1}{2}\sin(2\Omega t)(c_{1}\cos^{2}(\Omega t)-c_{2}\sin^{2}(\Omega t)).$
Using the identities
$\cos^{4}(\Omega t)+\sin^{4}(\Omega t)=1-\frac{1}{2}\sin^{2}(2\Omega t)$
and
$c_{1}\cos^{2}(\Omega t)-c_{2}\sin^{2}(\Omega t)=\frac{1}{2}(c_{1}-c_{2})+(c_{1}+c_{2})\cos(2\Omega t),$
we obtain the following estimates for the coefficients of system (8): for all $t\geq 0$,
$k_{11}\leq$ $\kappa_{1}(t)\leq k_{12},$ (9)
$|$ $\kappa_{2}(t)|\leq k_{2},$
where $k_{11}=\frac{1}{2}\min\{c_{1},c_{2}\}$, $k_{12}=\max\{c_{1},c_{2}\}$, $k_{2}=\frac{1}{4}|c_{1}-c_{2}|+\frac{1}{8}(c_{1}+c_{2})$. To prove the exponential stability of the trivial solution of system (8), consider the function
$V(\xi,\eta)=\frac{1}{2}\xi^{2}+\frac{1}{2}\eta^{2}+\gamma\xi\eta$ (10)
with $\gamma\in(0,1)$ to be defined. Then
$\dot{V}=$ $-(\kappa_{1}(t)-\gamma\Omega)\xi^{2}-\gamma(\Omega+\kappa_{2}(t))\eta^{2}$
$-(\gamma\kappa_{1}(t)+\kappa_{2}(t))\xi\eta$
$\leq$ $-\alpha_{1}\xi^{2}-\alpha_{2}\eta^{2}+\alpha_{12}|\xi\eta|$
with
$\alpha_{1}=(k_{11}-\gamma\Omega),\,\alpha_{2}=\gamma(\Omega+k_{2}),\,\alpha_{12}=(\gamma k_{12}+k_{2}).$
Requiring $\gamma<\frac{k_{11}}{\Omega}$, we ensure $\alpha_{1}>0$. Thus, to have negative definetness of the function $\dot{V}$ it is enough to ensure
$\alpha_{12}^{2}-4\alpha_{1}\alpha_{2}<0,$
which is equivalent to the requirement
$(4\Omega^{2}+4k_{2}\Omega+k_{12}^{2})\gamma^{2}-2(2\Omega k_{11}+2k_{11}k_{2}-k_{2}k_{12})\gamma+k_{2}^{2}<0.$ (11)
To ensure that the latter inequality is solvable, it is enough to have
$2(2\Omega k_{11}+2k_{11}k_{2}-k_{2}k_{12})^{2}-4k_{2}^{2}(4\Omega^{2}+4k_{2}\Omega+k_{12}^{2})\geq 0.$
Factorizing the left hand side, we obtain
$16(\Omega+k_{2})(\Omega(k_{11}^{2}-k_{2}^{2})+k_{11}k_{2}(k_{11}-k_{12}))\geq 0,$
which, in turn, leads to the requirement
$\Omega(k_{11}^{2}-k_{2}^{2})+k_{11}k_{2}(k_{11}-k_{12})\geq 0.$ (12)
Because of the definition of $k_{11},k_{12}$, the difference $k_{11}-k_{12}$ is always negative, while by the conditions of the Theorem, $k_{11}^{2}-k_{2}^{2}>0$ and
$\Omega\geq\frac{k_{11}k_{2}(k_{12}-k_{11})}{k_{11}^{2}-k_{2}^{2}}.$
Thus, inequality (12) is satisfied, which means that there exists a $\hat{\gamma}>0$ such that, for all $\gamma\in(0,\hat{\gamma}$], requirement (11) is satisfied. Thus, if $\gamma\in(0,\min\{1,\hat{\gamma},\frac{k_{11}}{\Omega}\}]$, then $\alpha_{1}\xi^{2}-\alpha_{2}\eta^{2}+\alpha_{12}|\xi\eta|$ is negative definite, therefore, such that
$\dot{V}\leq-\mu(\xi^{2}+\eta^{2}),$
where $\mu>0$ is the greatest eigenvalue of the matrix $\left(\begin{matrix}\alpha_{1}&\alpha_{12}/2\\ \alpha_{12}/2&\alpha_{2}\end{matrix}\right)$. Similarly,
$\frac{1-\gamma}{2}(\xi^{2}+\eta^{2})\leq V\leq\frac{1+\gamma}{2}(\xi^{2}+\eta^{2}).$
Thus,
$\dot{V}\leq-\frac{2\mu}{1+\gamma}V,$
which yields the exponential decay
$V(t)\leq V(0)e^{-\frac{2\mu t}{1+\gamma}},$
and
$\xi^{2}+\eta^{2}\leq\frac{1+\gamma}{1-\gamma}(\xi(0)^{2}+\eta(0)^{2})e^{-\frac{2\mu t}{1+\gamma}}.$
Coming back to the $(\bar{x},\bar{y})$-variables, we conclude
$(\bar{x}-\bar{x}_{d})^{2}+(\bar{y}-\bar{y}_{d})^{2}\leq\frac{1+\gamma}{1-\gamma}(\xi(0)^{2}+\eta(0)^{2})e^{-\frac{2\mu t}{1+\gamma}}.$
Remark 3.1
The simplest case in which the conditions of Theorem 1 are satisfied is when $c_{1}=c_{2}=c>0$, $\Omega>\frac{c}{3}$. We emphasize that the conditions on $C_{1},C_{2},\Omega$ are sufficient but not necessary, as they result from the particular choice of the Lyapunov function used in the proof. We expect that more general conditions could be derived by using, for example, a Lyapunov function with time-periodic coefficients or by applying Barbalat’s lemma. We leave this question for future work.
Theorem 2
The unicycle system (6) with $h=0$ is practically exponentially stable for any compact set $D\subset\mathbb{R}^{2}$ such that $(x_{d},y_{d})\in D$.
Proof. Let us define the vectors $\bm{X}=(x,y)\in\mathbb{R}^{2}$, $\bm{X}_{d}=(x_{d},y_{d})\in\mathbb{R}^{2}$ and $\bm{\bar{X}}=(\bar{x},\bar{y})\in\mathbb{R}^{2}$. We have:
[…]
For any initial condition $\bm{X}_{0}=\bm{X}(0)=\bm{\bar{X}}(0)$, from Theorem 1, there exists $q_{1}>0$ and $q_{2}>0$ such that $\bm{\bar{X}}-\bm{X}_{d}||\leq q_{1}e^{-q_{2}t}$. Moreover, per (Pokhrel and Eisa, 2023, Theorem 4), there exists $d(\varepsilon)\to 0$ as $\varepsilon\to 0$ such that $\bm{X}-\bm{\bar{X}}||\leq d(\varepsilon)$ for all $T>0$ and $t\in[0,T]$. Hence, the inequality (13) becomes:
$\bm{X}-\bm{X}_{d}||\leq q_{1}e^{-q_{2}t}+d(\varepsilon),\>\>\forall T>0\>\text{and}\>t\in[0,T].$ (14)
Remark 3.2
The inequality (14) guarantees that for any positive time $T$, which can be made large as needed, the trajectories of the unicycle system (6) will decay exponentially to a neighborhood about the extremum $\bm{X}_{d}$. Said neighborhood can be made arbitrarily small via the parameter $\varepsilon$. The reader can refer to Khalil and Grizzle (2002); Maggia et al. (2020); Pokhrel and Eisa (2023) for more details on the concept of practical exponential stability.
4 Simulation and Experimental Results
In this section, we present the simulation results, the experimental setup, and the experimental validation for the proposed exponentially convergent unicycle design. To start, we briefly review the experimental setup, as shown in Figure 2.
4.1 Experimental Setup
For experiments where the objective function is known, a motion capture system (MCS) is used to track the robot’s position in the $(x,y)$ coordinates. For the light source-seeking experiment, an analog light sensor connected to an Arduino Nano ESP32 board is mounted on the robot to measure light intensity. The robot used in our experiment is TurtleBot3 (see detailed information about this robot in Bajpai (2024); Bajpai et al. (2024)). The sensor readings are transmitted to the computer and used as feedback for the ESC system. The MCS is also used in this case to record and observe the robot’s position for performance evaluation. The light sensor consists of a small photoresistor that measures the intensity of light incident on its surface. As the light intensity increases, the sensor output decreases; hence, minimizing the sensor’s measured value corresponds to approaching the point of maximum light intensity.
Figure 2: Modeling, Dynamics, and Control Lab (MDCL (2025)): (1) light source; (2) TurtleBot3 robot; (3) motion capture system (MCS); (4) analog light sensor; (5) Arduino Nano ESP32; (6) MCS markers.
4.2 Simulation Results
In this subsection, we compare the traditional ESC design commonly found in the literature that is based on first-order Lie bracket design (Dürr et al. (2013)) with the proposed exponentially convergent unicycle ESC design developed for a fourth-order objective function based on third-order Lie bracket. The objective function used in this study is defined as
$J(x,y)=(x-1)^{4}+(y+2)^{4},$
which attains its minimum at $(x_{d},y_{d})=(1,-2)$. We remark here that this objective function satisfies the condition required for Theorem 1 (see Remark 1). The optional high-pass filter (HPF) is disabled in these simulations to isolate the effect of the proposed control law (i.e., $h=0$). The complete set of simulation parameters is provided in Table 1, and identical values are used for both ESC designs to ensure a fair comparison.
Table 1: Simulation Parameters
Parameter Value
$C_{1},\ C_{2}$ 1, 11,\ 1
$a$ $0.5$
$c$ $0.5$
$\varepsilon$ $0.001$
$\Omega$ 1.4rad/s1.4\ $\text${rad/s}
x(0),y(0)x(0),\ y(0) 1.6,1.4m1.6,\ -1.4\ $\text${m}
$x_{d},\ y_{d}$ 1,2m1,\ -2\ $\text${m}
Figure 3 illustrates the time histories of $x$ and $y$ as well as the planar trajectory of the unicycle. The results clearly demonstrate that the proposed exponentially convergent ESC achieves significantly faster convergence toward the desired equilibrium compared with the traditional ESC method from literature based on first-order Lie bracket design (Dürr et al. (2013)). Specifically, for the fourth-order objective function considered, the traditional ESC fails to converge within the 100 s simulation window, whereas the proposed design successfully converges within approximately 20 s.
These results validate the theoretical predictions derived in Section III and highlight the capability of the proposed approach to handle higher-order objective functions with markedly improved transient response and convergence speed.
Figure 3: Simulation results comparing the proposed exponentially convergent unicycle ESC with the traditional ESC based on first-order Lie bracket design from Dürr et al. (2013) for a fourth-order objective function: (a) $x$ position, (b) $y$ position, and (c) planar trajectory of the proposed unicycle design.
4.3 Experimental Results
We now extend the validation of the proposed exponentially convergent unicycle ESC design to real-world experiments. The tests are conducted using the same fourth-order objective function employed in the simulation study, with the addition of the high-pass filter (HPF) to improve transient performance. The complete set of parameters for both the traditional and exponentially convergent unicycle ESC designs is listed in Table 2.
Table 2: Experimental Parameters for Both Traditional and Exponentially Convergent Unicycle ESC Designs with a Fourth-Order Objective Function
Parameter Value
$C_{1},\ C_{2}$ 1, 11,\ 1
$a$ $0.01121$
$c$ $10$
$\varepsilon$ $0.2992$
$\Omega$ 1.4rad/s1.4\ $\text${rad/s}
$e$ 11
x(0),y(0)x(0),\ y(0) 1.6,1.4m1.6,\ -1.4\ $\text${m}
$x_{d},\ y_{d}$ 1,2m1,\ -2\ $\text${m}
We note here that our design parameters meet the condition of Theorem 1 (see Remark 1). The experimental results are presented in Figures 4 and 5. As shown in Figure 4, the traditional ESC method based on first-order Lie bracket from Dürr et al. (2013) fails to converge to the true minimum of the objective function even after 1200 s, instead settling near (x,y)=(0.98,1.93)(x,y)=(0.98,-1.93).
In contrast, the exponentially convergent unicycle ESC design converges much faster, first reaching the minimum at approximately 350 s and remaining about the true extremum for the duration of the experiment. The new ESC design clearly outperforms the classic literature ESC design and is validated by its ability to reach the minimum in significantly less time than the traditional approach. It is also important to note that a fourth-order objective function exhibits relatively flat regions near the minimum compared to a second-order objective function. The exponentially convergent unicycle ESC design can successfully steer the robot through these flat regions, as evident in Figure 6. Moreover, the variations in the objective function values become very small when the robot is close to the extremum point, confirming that steady-state convergence has been achieved. It is also important to highlight that the proposed design performed well even with expected resolution issues from the motion capturing system feedback due to the very small values of the fourth-order objective function, especially near the extremum. This is a positive indication about the robustness of the proposed design. The reader is directed to watch the experiment in our YouTube channel (Palanikumar et al. (2025a)).
Figure 4: Experimental results for the traditional ESC design based on first-order Lie bracket from Dürr et al. (2013) with a fourth-order objective function. (a) $x$ position (b) $y$ position (c) planar trajectory of the robot. Note that the x-position converges to $0.98$ meters (black-dashed), resulting in a 0.02 meter error, and the y-position converges to 1.93-1.93 (black-dashed), which results in a 0.07 meter error. The true minimum is in red.
Figure 5: Experimental results for the exponentially convergent unicycle ESC design with a fourth-order objective function. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (Palanikumar et al. (2025a)).
Figure 6: Planar trajectory of the robot with exponentially convergent unicycle ESC design plotted on the objective function distribution. The robot successfully navigates through relatively flat regions where the objective function values are extremely small.
For further verification and validation, we conducted an additional experiment to demonstrate the model-free nature of the proposed exponentially convergent unicycle ESC design. In this test, a light source-seeking task was performed in which the measured light intensity was directly used as feedback to the ESC controller. A high-pass filter (HPF) was again employed to mitigate the influence of natural fluctuations in the light sensor readings. The complete set of parameters used in this experiment is listed in Table 3. The point of maximum light intensity corresponds to the position where the photoresistive element of the sensor receives the highest illumination, which was determined experimentally to be located at (x,y)=(0.8035,2.202)(x,y)=(0.8035,-2.202).
Table 3: Experimental Parameters for Light Source Seeking with Exponentially Convergent Unicycle ESC Design
Parameter Value
$a$ $0.006665$
$c$ $0.001$
$\varepsilon$ $0.1496$
$\Omega$ 1.4rad/s1.4\ $\text${rad/s}
$e$ 66
x(0),y(0)x(0),\ y(0) 1.3,1.7m1.3,\ -1.7\ $\text${m}
The results of this experiment are presented in Figure 7. As shown, the robot successfully reaches and oscillates around the location of maximum light intensity. The light sensor measurements decrease as the robot approaches the source, and the mean value of the signal is minimized toward the end of the experiment, despite small oscillations caused by the robot’s continuous motion. These results further confirm that the proposed unicycle ESC design can autonomously steer the robot through an unknown and spatially varying light field, thereby validating its real-time, model-free source-seeking capability. The reader is directed to watch this experiment video in YouTube (Palanikumar et al. (2025b)).
Figure 7: Experimental results for light source-seeking with the exponentially convergent ESC model. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (Palanikumar et al. (2025b)).
5 Conclusion and future work
This paper provided a novel, first-of-its-kind exponentially convergent unicycle design inspired by significant recent results in higher-order Lie bracket approximations (Grushkovskaya and Eisa (2025); Pokhrel and Eisa (2023)). We proved exponential stability of the proposed design for objective functions that behave locally as fourth-degree polynomial. Additionally, we validated the results by simulations, and more importantly, via experiments. In our experiments, we did not only validate the ability of the proposed design to operate with known objective functions, but we also validated its ability to operate in a complete model-free condition in a light source-seeking experiment.
In the future, we aim at generalizing the results here in this paper to general higher-degree polynomial objective functions and expand the proposed design into different forms that may reduce oscillations and posses bounded update rate.